OpenStudy (ttforver):

List the steps -12/3*(-8(-4)^2-6)+2

OpenStudy (welshfella):

have you heard of PEMDAS - the order of operations.

OpenStudy (ttforver):

yes i have

OpenStudy (ttforver):

P-Parenthesis E-Exponent M-Multiply D-Divide A-Adding S-Subtracting

ILovePuppiesLol (ilovepuppieslol):

now list the steps of the equations in that order and ur done

alones (alones):

So this how your math problem looks like? \[\frac{ -12 }{ 3 }(-8(-4^2)-6) +2\]

OpenStudy (ttforver):

yes

alones (alones):

Okay so let's divide 012 by 3...

alones (alones):

Oh i mean -12

OpenStudy (ttforver):

:)

ILovePuppiesLol (ilovepuppieslol):

we have to do the parentheses first

OpenStudy (ttforver):

ok

ILovePuppiesLol (ilovepuppieslol):

what is the equation inside the parentheses?

alones (alones):

\(\frac{ -12 }{ 3 }\)=?

OpenStudy (ttforver):

4

alones (alones):

do the perenthesis first

alones (alones):

*parentheses

ILovePuppiesLol (ilovepuppieslol):

do u have a screenshot of ur question? because the way u wrote it is kinda confusing

OpenStudy (ttforver):

yeah i do hold on

OpenStudy (ttforver):

how do i post a screenshot

ILovePuppiesLol (ilovepuppieslol):

attach file, or if ur using lightshot, juust post the link

OpenStudy (ttforver):

dang its not doing nothing

OpenStudy (ttforver):

do you want me to try to type it again

ILovePuppiesLol (ilovepuppieslol):

its fine, im kinda busy rn, try tagging others, sorry

ILovePuppiesLol (ilovepuppieslol):

just use PEMDAS, and order the steps

OpenStudy (welshfella):

OK after parentheses comes the exponent can you simplify the exponent part ?

OpenStudy (welshfella):

(-4)^2 = ?